OpenStudy (anonymous):

what is the derivative of 3^-x

OpenStudy (anonymous):

\[3^{-x}=\frac{ 1 }{ 3^{x}}\] now you can use the quotient rule.\[\frac{ 0(3^{x})-3^{x} \ln 3 }{ 3^{2x} }\] remember the derivative of the constant is zero so 0(3^x) cancels out. now simplifying it would look like \[-\frac{ \ln 3 }{ 3^{x} }\]